# Semantic Residual Prompts for Continual Learning

### v1

@@ Abstract
Prompt-tuning methods for Continual Learning (CL) freeze a large pre-trained model and focus training on a few parameter vectors termed prompts. Most of these methods organize these vectors in a pool of key-value pairs, and use the input image as query to retrieve the prompts (values). However, as keys are learned while tasks progress, the prompting selection strategy is itself subject to catastrophic forgetting, an issue often overlooked by existing approaches. For instance, prompts introduced to accommodate new tasks might end up interfering with previously learned prompts.
To make the selection strategy more stable, we ask a foundational model (CLIP) to select our prompt within a two-level adaptation mechanism. Specifically, the first level leverages standard textual prompts for the CLIP textual encoder, leading to stable class prototypes. The second level, instead, uses these prototypes along with the query image as keys to index a second pool. The retrieved prompts serve to adapt a pre-trained ViT, granting plasticity. In doing so, we also propose a novel residual mechanism to transfer CLIP semantics to the ViT layers.
Through extensive analysis on established CL benchmarks, we show that our method significantly outperforms both state-of-the-art CL approaches and the zero-shot CLIP test. Notably, our findings hold true even for datasets with a substantial domain gap w.r.t. the pre-training knowledge of the backbone model, as showcased by experiments on satellite imagery and medical datasets.
@@ 1 Introduction
The emergence of large-scale pre-trained Transformers [41] and foundation models [1] has recently changed the adaptation paradigm, usually moving towards parameter-efficient plasticity, in which only a few parameters are optimized to solve a new task while keeping most of the model frozen. A successful example is prompt tuning, in which a few learnable vectors (“soft prompts”) are appended to the input image tokens. For instance, CoOp [50] learns a context prompt, which is concatenated with the textual name of the class label and fed to the textual encoder of CLIP [32]. The latter generates a class prototype in feature space, which is used to classify the images of the downstream task. Similarly, VPT [17] learns task-specific visual prompts for each layer of an ImageNet pre-trained Vision Transformer (ViT) [10], while CoCoOp [49] extends CoOp by conditioning its prompts on the input image.
@@
Prompt tuning has recently been adopted in different CL proposals [47, 46, 40, 43, 5] with good results. These methods, inspired by CoOp, CoCoOp, and VPT, usually devise a shared pool of key-value pairs to manage the prompts learned during each round. As shown in Fig. 1 (left), the input image is used as a query to retrieve a subset of relevant prompts from the pool. To make the selection strategy effective, the keys are learned along with the values (i.e., the prompts). The learning criteria for these keys usually involve a form of weak supervision that pulls selected keys closer to corresponding query features, with further orthogonality constraints to encourage diversification.
@@
Despite the extensive application in incremental scenarios, prompting approaches fall into a pitfall, which is the main focus of our work. As the key space is continuously updated with no further access to past queries/tasks, the selection mechanism is itself subject to catastrophic forgetting and possible misalignments between queries and keys. As outlined in Sec. 3.4, this causes interference in the prompt selection (i.e., examples of a certain task are mapped to unrelated tasks), and hence to sub-optimal performance.
@@
We hence propose to employ a foundation model like CLIP to realize the query-key matching mechanism, yielding a more stable selection strategy (see Sec. 3.4). In fact, it has been shown that the fine-tuning regime of CLIP models favors stability over plasticity [5], as distinct parameters can be naturally devoted to distinct concepts, mitigating their interference. However, to enhance plasticity, we use the prompts retrieved through CLIP to condition another backbone, specifically an ImageNet pre-trained ViT. Notably, this leads to a two-level prompting pipeline (Fig. 1 right), which is the main technical contribution of our work. Briefly: as in CoOp and AttriCLIP [44], we train class-specific prompts (first level prompts) to be fed to the CLIP textual encoder. Conversely, we exploit the resulting embedding not for computing classification scores, but as keys of a second prompt pool. The retrieved second-level prompts are used to adapt the ImageNet pre-trained ViT, providing the final output prediction.
@@
Our secondary contribution regards the mechanism introduced to prompt the pre-trained ViT. We indeed propose an additive residual mechanism to adapt the MLPs of a Transformer. While most methods [46, 40] concatenate the learned prompts to the arguments of the multi-head self-attention (MSA) layer (Sec. 2), our second-level prompts are used as semantic residuals which are added before the MLP layer input. While there is no guarantee that a pre-trained and frozen MSA layer will take into account all the prompt tokens, our additive mechanism forces the Transformer embeddings to include CLIP-derived semantics.
@@
Finally, inspired by Slow Learner with Classifier Alignment (SLCA) [48] [48], where first- and second- order feature statistics are used to generate synthetic samples of past tasks (Sec. 2), we adopt a similar generative replay method in both stages of our approach. However, differently from SLCA, we use a Mixture of Gaussians (MoGs) to model the multi-modal distribution of each class, leading to a more effective generation process.
@@
To assess our intuitions, we test our approach termed STAR-Prompt (Semantic Two-level Additive Residual Prompt) on nine image classification datasets. These encompass well-established CL benchmarks, such as Split Imagenet-R and Split CIFAR-100 [46, 40, 48, 5], but also satellite and medical settings, hence covering domains that differ significantly  [28, 8, 32] from the pre-training source domains of both ImageNet and CLIP. As our experiments outline, prompting-based CL approaches struggle with such large domain gaps, often being outperformed by older methods relying on the good old schema of rehearsal.
The results indicate that we significantly outperform the state-of-the-art, getting a remarkable $+5.84$ improvement of final accuracy, with peaks of $+3$ on Split Imagenet-R and $+17$ on Split Cars-196. Importantly, we prove the resilience of our framework to severe domain shifts. For example, when the zero-shot CLIP model struggles (e.g., achieving only $57.74$ accuracy on Split EuroSAT), our two-level approach instead excels (i.e., $94.24$), underscoring the merits of a stable and continuous adaptation over the simple reuse of frozen large pre-trained models.
We therefore acknowledge the following contributions:
@@
  • We shed light on the stability issue of prompt selection, and propose a two-level prompting strategy that builds upon the stability of foundation models.
  • A novel prompting approach based on semantic residuals is presented.
@@
  • We deepen the current understanding of incremental adaption, underscoring that large pre-trained models still requires adaptability.
@@ 2 Related work
Continual Learning. CL approaches are usually classified according to the strategy they adopt against forgetting. Regularization-based methods adjust the loss function to restrict significant changes in either the model weights or its activations when moving to other tasks [20, 23, 35, 27]. Architectural methods alter the structure of the model by devoting separate groups of parameters to different tasks [37, 25, 38]. Finally, Rehearsal-based methods alleviate forgetting by storing a subset of samples from past tasks in a memory buffer, which are then replayed along with the data of the current task [33, 34, 14, 2, 3]. A sub-category of rehearsal methods is based on generative replay, where real samples are replaced by synthetically generated samples [36, 39, 18].
Prompt tuning methods. The first attempt to use prompt tuning in CL is Learning to Prompt (L2P[47], where a pool of prompts is shared by all tasks, and the input image is used as the query to select the most appropriate prompts from this pool. DualPrompt [46] introduces a hierarchy of general and task-specific prompts and adopts prefix-tuning over prompt tuning, where the former prepends prompts to the keys and the values of an MSA layer rather than prepending prompts to the input tokens. Prefix-tuning is also adopted in CODA-Prompt [40], which additionally introduces a query weighting mechanism and an orthogonality soft-constraint to encourage prompt independence (both of which we also adopt, see Sec. 3.3).
Recently, several works have investigated the application of the CLIP model [32] in incremental settings. In AttriCLIP [43] the prompts are fed to the CLIP’s textual encoder, while PromptFusion [5] takes inspiration from VPT and trains prompts for the CLIP’s textual encoder jointly with visual prompts for a pre-trained ViT. Differently from PromptFusion, which uses these prompts in parallel fusing the corresponding classification scores, we instead employ them sequentially. Specifically, the textual prompts are used to produce class-specific keys to index a second pool of prompts.
Figure 2: The architecture of STAR-Prompt. The bottom left box illustrates our CL setting, in which first- and second-level prompts of the tasks are frozen.
@@ 3 Method
Problem setting. In CL, a model is trained on a sequence of $T$ tasks, where each task $D_{t}$, $t=1,\dots,T$ is composed of sample-label pairs $(x_{i},y_{i})$, $i=1,\dots,|D_{t}|$. In this paper, we adopt a class-incremental setting, in which the set of classes ${\cal Y}_{t}$ of task $D_{t}$ is disjoint from ${\cal Y}_{t^{\prime}}$ (with $t\neq t^{\prime}$), and the task identity ($t$) is unknown at inference time. For simplicity, we assume that $|{\cal Y}_{1}|=\dots=|{\cal Y}_{T}|=N$. Importantly, when the model is trained on $D_{t}$, the training data from the old tasks $D_{1:t-1}$ are inaccessible.
@@ 3.1 Overview
In a nutshell, every parameter of these networks is frozen during training. To adapt them in a parameter-efficient manner while minimizing interference, we devise a two-level codebook strategy. The first one provides learnable context passed as input to the CLIP text encoder: by doing so, we produce stable class prototypes. The second level treats these textual prototypes as keys of a second codebook of prompts, which serve to provide a more plastic behaviour to the second model $f(\cdot)$.
Specifically, the first-level prompting stage resembles CoOp and conditions the text encoder $E_{txt}(\cdot)$ by concatenating a class-specific learnable vector to the input sequence. However, differently from CoOp, we do not directly exploit the learned textual class-prototypes to perform classification. In fact, we use them as keys for comparison against the visual query, derived by applying the CLIP visual encoder $E_{vis}(\cdot)$ on the input image. Based on the resulting similarities, we retrieve a second-level learnable prompt R, termed semantic residual vector, aiming to transfer class-specific semantics to the ViT $f(\cdot)$.
@@ 3.2 STAR-Prompt: Two-Level Prompt Tuning
In the first prompting stage, we associate a learnable prompt $\boldsymbol{p}_{c}$ to each class $y_{c}\in{\cal Y}_{t}$. The prompt $\boldsymbol{p}_{c}$ consists of a single learnable token that is concatenated with the initial word embedding of the class $y_{c}$, indicated with $\operatorname{[CL-NAME]}$:
$\boldsymbol{w}_{c}=E_{txt}([\boldsymbol{p}_{c};\operatorname{[CL-NAME]}]).$ (4)
For instance, if $y_{c}$ refers to the class ‘dog’, then $\operatorname{[CL-NAME]}$ is the embedding of the word ‘dog’. These two embeddings are fed to $E_{txt}(\cdot)$ to get a final textual representation $\boldsymbol{w}_{c}\in\mathbb{R}^{d}$.
While CoOp utilizes $\boldsymbol{w}_{c}$ as a class prototype for image classification of $x$, we treat $\boldsymbol{w}_{c}$ as a class-prototype key. Specifically, we use these class-specific keys to retrieve a second set of learnable prompts $\mathbf{Q_{c}}$. These prompts aim to adapt each layer of the other architecture, i.e., the ImageNet pre-trained Vision Transformer $f(\cdot)$. In particular, let $d^{\prime}$ and $L$ denote the hidden dimension and the number of layers of $f(\cdot)$ respectively. Then, $Q_{c}[l]\in\mathbb{R}^{d^{\prime}}$ represents the second set of class-specific learnable prompts used to adapt the $l$-th layer of $f(\cdot)$.
In details, the second-level prompting stage leverages the CLIP multi-modal embedding space. Given the visual embedding of the input image $\boldsymbol{z}=E_{vis}(x)$ (query), we compute its cosine similarity $\operatorname{sim}_{c}=\langle\boldsymbol{z},\boldsymbol{w}_{c}\rangle$ with each class prototype key $\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{Nt}$. Each of these keys is derived from the first-level prompting stage, and corresponds to one of the $Nt$ classes observed so far.
Afterwards, we build the semantic residual $\mathbf{R}$ by using the prompt $Q_{c_{k}}$ of the class with the highest cosine similarity, as follow:
$R$ $=\operatorname{sim}_{c_{k}}\cdot\ Q_{c_{k}},\;$ (5)
$\textrm{where}\quad c_{k}$ $=\operatorname*{arg\,max}_{c=1,\dots,Nt}\operatorname{sim}_{c_{k}}.$ (6)
@@
The residuals $R$ are then used to transfer CLIP-based class-specific cues to each MLP layer of $f(x)$. We hence replace Eq. 1 with:
$\boldsymbol{e}^{\prime}=\operatorname{MSA}(\operatorname{LN}(\boldsymbol{e}_{l}))+\boldsymbol{e}_{l}+R[l].$ (7)
Main training loss. For each task $D_{t}$, we extend the two codebooks adding the class-specific prompts of the current task. In detail, the set of first-level prompts is computed as: ${\cal P}_{t}:={\cal P}_{t-1}\cup\{\boldsymbol{p}_{c}|c\in{\cal Y}_{t}\}$, while the second-level codebook is obtained by: ${\cal Q}_{t}:={\cal Q}_{t-1}\cup\{Q_{c}|c\in{\cal Y}_{t}\}$. To avoid forgetting, we freeze the prompts of the old tasks (${\cal P}_{t-1}$ and ${\cal Q}_{t-1}$) and train only those corresponding to the current task. Although both levels may be trained jointly and end-to-end, we strive for simplicity and structure the optimization process as a two-stage training procedure. Specifically, for both codebooks, we use the cross-entropy as the main loss function:
@@
where, in the first stage, the posteriors are computed as:
@@
where $\tau$ indicates the temperature parameter of the CLIP model. Conversely, in the second stage, the posteriors in Eq. 8 consider the $\operatorname{CLS}$ token $\boldsymbol{e}_{L}^{\operatorname{CLS}}=f(x_{i};R)$ of the last layer of $f(\cdot)$ [10], as follows:
@@
where $g_{\theta_{t}}(\cdot)$ represents the classification head of the current task $t$. This module comprises of a linear projection followed by the softmax layer. The parameters $\theta_{t}$ of the projection layer are initialized at the beginning of task $t$, and trained while keeping the weights of the old tasks frozen.

### v3

@@ Abstract
Prompt-tuning methods for Continual Learning (CL) freeze a large pre-trained model and train a few parameter vectors termed prompts. Most of these methods organize these vectors in a pool of key-value pairs and use the input image as query to retrieve the prompts (values). However, as keys are learned while tasks progress, the prompting selection strategy is itself subject to catastrophic forgetting, an issue often overlooked by existing approaches. For instance, prompts introduced to accommodate new tasks might end up interfering with previously learned prompts. To make the selection strategy more stable, we leverage a foundation model (CLIP) to select our prompts within a two-level adaptation mechanism. Specifically, the first level leverages a standard textual prompt pool for the CLIP textual encoder, leading to stable class prototypes. The second level, instead, uses these prototypes along with the query image as keys to index a second pool. The retrieved prompts serve to adapt a pre-trained ViT, granting plasticity. In doing so, we also propose a novel residual mechanism to transfer CLIP semantics to the ViT layers. Through extensive analysis on established CL benchmarks, we show that our method significantly outperforms both state-of-the-art CL approaches and the zero-shot CLIP test. Notably, our findings hold true even for datasets with a substantial domain gap w.r.t. the pre-training knowledge of the backbone model, as showcased by experiments on satellite imagery and medical datasets. The codebase is available at https://github.com/aimagelab/mammoth.
Keywords:
Continual learning Rehearsal-free Prompt-based learning
@@ 1 Introduction
The emergence of large-scale pre-trained Transformers [42] and foundation models [2] has recently changed the adaptation paradigm, usually moving towards parameter-efficient plasticity, in which only a few parameters are optimized to solve a new task while keeping most of the model frozen. A successful example is prompt tuning, in which a few learnable vectors (“soft prompts”) are appended to the input image tokens. For instance, CoOp [52] learns a context prompt, which is concatenated with the textual name of the class label and fed to the textual encoder of CLIP [33]. The latter generates a class prototype in feature space, which is used to classify the images of the task at hand. Similarly, Visual Prompt Tuning (VPT) [18] learns task-specific visual prompts for each layer of an ImageNet pre-trained Vision Transformer (ViT) [11], while CoCoOp [51] extends CoOp by conditioning its prompts on the input image.
@@
Prompt tuning has recently been adopted in different CL proposals [49, 48, 41, 45, 6] with good results. These methods, inspired by CoOp, CoCoOp, and VPT, usually devise a shared pool of key-value pairs to manage the prompts learned during each round. As shown in Fig. 1 (left), the input image is used as a query to retrieve a subset of relevant prompts from the pool. To make the selection strategy effective, the keys are learned along with the values (i.e., the prompts). The learning criteria for these keys usually involve a form of weak supervision that pulls selected keys closer to corresponding query features, with further orthogonality constraints to encourage diversification [41].
@@
Despite the extensive application in incremental scenarios, prompting approaches fall into a pitfall, which is the main focus of our work. As the key space is continuously updated with no further access to past queries/tasks, the selection mechanism is itself subject to catastrophic forgetting and possible misalignments between queries and keys. As outlined in Sec. 3.4, this causes interference in the prompt selection (i.e., prompts of a certain task are mapped to unrelated tasks), thus leading to sub-optimal performance.
@@
We hence propose to employ a foundation model like CLIP to realize the query-key matching mechanism, yielding a more stable selection strategy. In fact, it has been shown that the fine-tuning regime of CLIP models favors stability over plasticity [6], as distinct parameters can be naturally devoted to distinct concepts, mitigating their interference. However, to enhance plasticity, we use the prompts retrieved through CLIP to condition another backbone, specifically an ImageNet pre-trained ViT. Notably, this leads to a two-level prompting pipeline (Fig. 1 right), which is the main technical contribution of our work. Briefly: as in CoOp and AttriCLIP [46], we train class-specific prompts (first-level prompt pool) to be fed to the CLIP textual encoder. Conversely, we exploit the resulting embedding not for computing classification scores, but as keys of a second prompt pool. The retrieved second-level prompts are used to adapt the ImageNet pre-trained ViT, which provides the final output prediction.
@@
Our secondary contribution involves a novel mechanism to prompt the pre-trained ViT, utilizing an additive residual technique to adapt it. While most methods [48, 41] concatenate the learned prompts to the arguments of the Multi-head Self-Attention (MSA) layer (see Sec. 2), our second-level prompts are used as semantic residuals which are added before the MLP layer input. While there is no guarantee that a pre-trained and frozen MSA layer will take into account all the prompt tokens, our additive mechanism forces the ViT embeddings to include CLIP-derived semantics.
@@
Finally, inspired by SLCA [50], where first- and second-order feature statistics are used to generate synthetic samples of past tasks, we adopt a generative replay method in both stages of our approach. However, differently from SLCA, we use a Mixture of Gaussians (MoGs) to model the multimodal distribution of each class, leading to a more effective generation process.
@@
To assess our intuitions, we test our approach termed STAR-Prompt (Semantic Two-level Additive Residual Prompt) on nine image classification datasets. These encompass well-established CL benchmarks, such as Split Imagenet-R and Split CIFAR-100 [48, 41, 50, 6], but also satellite and medical settings, hence covering domains that differ significantly [29, 9, 33] from the pre-trainings of both ImageNet and CLIP. As our experiments outline, prompting-based CL approaches struggle with such large domain gaps, often being outperformed by older methods relying on the good old rehearsal schema. The results show a remarkable $\texttt{+}5.96$ average improvement in final accuracy w.r.t. the state of the art, with peaks of $\texttt{+}3$ on Split Imagenet-R and $\texttt{+}11$ on Split Cars-196. Importantly, we prove the resilience of our framework to severe domain shifts. For example, when the zero-shot CLIP model struggles (e.g., achieving only $54.50\%$ accuracy on Split EuroSAT), our two-level approach instead excels (i.e., $93.70\%$), underscoring the merits of a stable and continuous adaptation over the simple reuse of frozen large pre-trained models. We therefore acknowledge the following contributions:
@@
  • We shed light on the stability issue of prompt selection proposing a two-level prompting strategy that builds upon the stability of foundation models.
  • We present a novel prompting approach based on semantic residuals.
@@
  • We deepen the current understanding of incremental adaption, underscoring that large pre-trained models still require adaptability.
@@ 2 Related work
Continual Learning. CL approaches are usually classified according to the strategy they adopt against forgetting. Regularization-based methods adjust the loss function to restrict significant changes in either the model weights or its activations when moving to other tasks [21, 24, 36, 28]. Architectural methods alter the structure of the model by devoting separate groups of parameters to different tasks [38, 26, 39]. Finally, Rehearsal-based methods alleviate forgetting by storing a subset of samples from past tasks in a memory buffer, which are then optimized along with the data of the current task [34, 35, 15, 3, 4]. A sub-category of rehearsal methods is based on generative replay, where real samples are replaced by synthetically generated ones [37, 40, 19].
Prompting methods. The first attempt to use prompt tuning in CL is Learning to Prompt (L2P[49], where a pool of prompts is shared by all tasks, and the input image is used as the query to select the most appropriate prompts from this pool. DualPrompt [48] introduces a hierarchy of general and task-specific prompts and adopts prefix-tuning over prompt tuning, where the former prepends prompts to the keys and the values of MSA layers rather than prepending prompts to the input tokens. Prefix-tuning is also adopted in CODA-Prompt [41], which additionally introduces an end-to-end prompting mechanism and an orthogonality soft-constraint to encourage prompt independence (both of which we adopt, see Sec. 3.3).
Recently, several works have investigated the application of CLIP [33] in incremental settings. In AttriCLIP [45] the prompts are fed to the CLIP’s textual encoder, while PromptFusion [6] takes inspiration from VPT and trains prompts for the textual encoder jointly with visual prompts for a pre-trained ViT. Differently from PromptFusion, which uses these prompts in parallel fusing the corresponding classification scores, we instead employ them sequentially. Specifically, the textual prompts are used to produce class-specific keys to index a second pool of prompts.
Figure 2: The architecture of STAR-Prompt. The bottom left box illustrates our CL setting, in which first- and second-level prompts of the old tasks are frozen.
@@ 3 Method
Problem setting. In Continual Learning, a model is trained on a sequence of $T$ tasks, where each task $D_{t}$, $t=1,\dots,T$ is composed of sample-label pairs $(x_{i},y_{i})$, $i=1,\dots,|D_{t}|$. In this paper, we adopt a class-incremental setting [43], in which the set of classes ${\cal Y}_{t}$ of task $D_{t}$ is disjoint from ${\cal Y}_{t^{\prime}}$ (with $t\neq t^{\prime}$), and the task identity $t$ is unknown at inference time. For simplicity, we assume that $|{\cal Y}_{1}|=\dots=|{\cal Y}_{T}|=N$. Importantly, when the model is trained on $D_{t}$, the training data from the old tasks $D_{1:t-1}$ are inaccessible.
@@ 3.1 Overview
In a nutshell, every parameter of these networks is frozen during training. To adapt them in a parameter-efficient manner while minimizing interference, we devise a two-level codebook strategy. The first one provides learnable context passed as input to the CLIP text encoder: by doing so, we produce stable class prototypes. The second level leverages these text-level prototypes as keys of a second codebook of prompts, which serve to provide a more plastic behaviour to the second model $f(\cdot)$.
Specifically, the first-level prompting stage resembles CoOp and conditions the text encoder $E_{txt}(\cdot)$ by concatenating a class-specific learnable vector to the input sequence. However, differently from CoOp, we do not directly exploit the learned textual class prototypes to perform classification. In fact, we use them as keys for comparison against the visual query, derived by applying the CLIP visual encoder $E_{vis}(\cdot)$ on the input image. Based on the resulting similarities, we retrieve a second-level learnable prompt $R$, termed semantic residual vector, aiming to transfer class-specific semantics to the ViT $f(\cdot)$.
@@ 3.2 STAR-Prompt: Two-Level Prompt Tuning
In the first prompting stage, we associate a learnable prompt $p_{c}$ to each class $y_{c}\in{\cal Y}_{t}$. The prompt $p_{c}$ consists of a single learnable token that is concatenated with the initial word embedding of the class $y_{c}$, indicated with $\operatorname{[CLS-NAME]}$:
$\boldsymbol{w}_{c}=E_{txt}([p_{c};\operatorname{[CLS-NAME]}]).$ (4)
For instance, if $y_{c}$ refers to the class ‘dog’, then $\operatorname{[CLS-NAME]}$ is the embedding of the word ‘dog’. These two embeddings are fed to $E_{txt}(\cdot)$ to get a final textual representation $\boldsymbol{w}_{c}\in\mathbb{R}^{d}$.
While CoOp utilizes $\boldsymbol{w}_{c}$ as a class prototype to classify the image $x$, we treat $\boldsymbol{w}_{c}$ as a class-specific key. Specifically, we use these class-specific keys to retrieve a second set of learnable prompts $Q_{c}\in\mathbb{R}^{L\times d^{\prime}}$. These prompts aim to adapt each layer of the ImageNet pre-trained Vision Transformer $f(\cdot)$. In particular, let $l<L$ denote the $l$-th layer of $f(\cdot)$ and $d^{\prime}$ its embedding dimension. Then, $Q_{c}[l]\in\mathbb{R}^{d^{\prime}}$ represents the second-level class-specific prompts used to adapt the $l$-th layer of $f(\cdot)$.
In detail, the second-level prompting stage leverages the CLIP multi-modal embedding space. Given the visual embedding of the input image $\boldsymbol{z}=E_{vis}(x)$ (query), we compute its cosine similarity $\operatorname{sim}_{c}=\langle\boldsymbol{z},\boldsymbol{w}_{c}\rangle$ with each class prototype key $\boldsymbol{w}_{1},\dots,\boldsymbol{w}_{Nt}$. Each of these keys is derived from the first-level prompting stage and corresponds to one of the $Nt$ classes observed so far.
Afterwards, we build the semantic residual $\mathbf{R}\in\mathbb{R}^{L\times d^{\prime}}$ by using the prompt $Q_{c_{k}}$ of the class with the highest cosine similarity, as follow:
$\mathbf{R}$ $=\operatorname{sim}_{c_{k}}\cdot\ Q_{c_{k}},\;$ (5)
$\textrm{where}\quad c_{k}$ $=\operatorname*{arg\,max}_{c=1,\dots,Nt}\operatorname{sim}_{c}.$ (6)
@@
The residuals $\mathbf{R}$ are then used to transfer CLIP-based class-specific cues to each MLP layer of $f(x)$. We hence replace Eq. 1 with:
$\boldsymbol{e}^{\prime}=\operatorname{MSA}(\operatorname{LN}(\boldsymbol{e}_{l}))+\boldsymbol{e}_{l}+\mathbf{R}[l].$ (7)
Please note that $\mathbf{R}[l]$ is a single $d^{\prime}$-dimensional vector that applies uniformly across the entire sequence of visual tokens. In other words, to ensure dimensional compatibility for summation, $\mathbf{R}[l]$ is repeated for each token in the sequence.
Main training objective. For each task $D_{t}$, we extend the two codebooks adding the class-specific prompts of the current task. In detail, the set of first-level prompts is computed as: ${\cal P}_{t}:={\cal P}_{t-1}\cup\{p_{c}~{}|~{}c\in{\cal Y}_{t}\}$, while the second-level codebook is obtained by: ${\cal Q}_{t}:={\cal Q}_{t-1}\cup\{Q_{c}~{}|~{}c\in{\cal Y}_{t}\}$. To avoid forgetting, we freeze the prompts of the old tasks (${\cal P}_{t-1}$ and ${\cal Q}_{t-1}$) and train only those corresponding to the current task. Although both levels may be trained jointly and end-to-end, we strive for simplicity and structure the optimization process as a two-stage training procedure. Specifically, for both codebooks, we use the cross-entropy as the main loss function:
@@
where, in the first stage, the posterior probabilities are computed as:
@@
with $\tau$ indicating the temperature parameter of CLIP. Conversely, in the second stage, the posteriors in Eq. 8 consider the $\operatorname{CLS}$ token $\boldsymbol{e}_{L}^{\operatorname{CLS}}=f(x_{i};\mathbf{R})$ of the last layer $L$ of $f(\cdot)$ [11], as follows:
@@
where $g_{\theta_{t}}(\cdot)$ represents the classification head of the current task $t$. This module comprises of a linear projection followed by the softmax layer. The parameters $\theta_{t}$ of the projection layer are initialized at the beginning of task $t$, and trained while keeping the classification heads of the old tasks frozen.